theresa sold magazine subscription with three prices: $18, $16, $25. She sold 1 fewer of the $18 subscriptions than of the $16 subscriptions and sold a total of 24 subscriptions. If her total sales amounted to $479,how many $25subscriptions did theresa sell?
Answer by lwsshak3(11628)

theresa sold magazine subscription with three prices: $18, $16, $25. She sold 1 fewer of the $18 subscriptions than of the $16 subscriptions and sold a total of 24 subscriptions. If her total sales amounted to $479,how many $25subscriptions did theresa sell?
***
let x=number of $16 subscriptions theresa sold
x-1=number of $18 subscriptions theresa sold
24-x-(x-1)=25-2x=number of $25 subscriptions theresa sold
..
16x+18(x-1)+25(25-2x=479
16x+18x-18+625-50x=479
16x=128
x=8
25-2x=25-16=9
number of $25 subscriptions theresa sold=9